Omar is 14 years old and his father is 35 years old.When will omar be four sevenths his fathers age?
Answer by dkppathak(439)

Omar is 14 years old and his father is 35 years old.When will omar be four sevenths his fathers age?
GIVEN THAT
present age of omar is 14 yrs
and father present age is 35 yrs
as per given conditions when omar will be 4/7 of father age
let it will be after X years
age of omar after X years = X+14
father age after X years = X+35
(x+14) = 4/7(x+35) as per given conditions
7(x+14)= 4(x+35)
7x+98 =4x +140
7x-3x=140-98
3x = 42
x=42/3= 14 years

Answer
omar will be 4/7th of father after 14 years